FOURFOLD CLUSTERS OF ROVIBRATIONAL ENERGIES IN H_{2}Te AND H_{2}Po

P. C. GÓMEZ, Dep. Química Física I, Facultad Química, U. Complutense de Madrid, E-28040 Madrid, Spain; L. F. PACIOS, Dep. Química y Bioquímica, E.T.S.I. Montes, U. Politécnica de Madrid, E-28040 Madrid, Spain; \underline{PER JENSEN}, FB 9 – Theoretische Chemie, Bergische Universität

– Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany.

An interesting novel pattern in rotation-vibration spectra is the cluster effect (i. e., the formation of four-member groups of nearly degenerate rotation-vibration energy levels at higher J and K_{a} values^{a}) in the molecules H_{2}S, H_{2}Se, H_{2}Te, and H_{2}Po. In the cluster states, local mode behaviour is induced by rotational excitation so that it occurs in the lowest vibrational states. We report here investigations of the cluster effect in the vibrational ground states of the H_{2}Te and H_{2}Po molecules. For H_{2}Te, we have obtained a new potential energy function by fitting to the available spectroscopic data, and for H_{2}Po, we have calculated the potential energy function ab initio by means of the CCSD(T) method, using an Averaged Relativistic Effective Potential for Po. With the new potential energy functions, the rotation-vibration energies of H_{2}^{130}Te and H_{2}^{209}Po have been calculated with the MORBID (Morse Oscillator Rigid Bender Internal Dynamics) Hamiltonian and computer program. In particular, we have calculated the rotational energy manifolds for J ≤ 40 in the vibrational ground states. An interesting aspect of the cluster effect is the change in vibrational ground state cluster formation taking place for the sequence of molecules H_{2}O → H_{2}S → H_{2}Se → H_{2}Te → H_{2}Po. The H_{2}O molecule does not exhibit any discernible cluster formation at J ≤ 40, but for all the other molecules the cluster size decreases rapidly with increasing J. For H_{2}Te and H_{2}Po, in which the central mass is effectively infinite compared with the hydrogen mass, a limiting behaviour is reached which would not change further if the central mass were made heavier. This work was supported by the European Commission through contract no. CHRX-CT94-0665 “Highly Excited Rovibrational States”. One of us (P. C. G.) would also like to acknowledge support from the Spanish DGICYT (Project PB94-1526).

aP. Jensen, G. Osmann, and I. N. Kozin, in: Advanced Series in Physical Chemistry: Vibration-Rotational Spectroscopy and Molecular Dynamics (D. Papoušek, Ed.), World Scientific Publishing Company, Singapore, in press, and references therein.